OpenStudy (anonymous):

Write a sequence that has two geometric means between 48 and -750 I think the sequence should be 48, -120, 300, -750 I got that sequence by multiplying each number by -2.5 I want to know if the answer I got is correct

OpenStudy (amistre64):

find the range and divide by 3

OpenStudy (amistre64):

essentially youre setting up an arithmetic sequence is another way to view it

OpenStudy (amistre64):

geometric ... i always read over that for some reason

OpenStudy (anonymous):

-750+48=-702/3=-234?

OpenStudy (amistre64):

that would be the difference for finding 2 arithmetic means; i read over the geometric part

OpenStudy (anonymous):

I don't understand what a geometric mean is? Does it mean to fill in the two spots between the two numbers to finish the sequence?

OpenStudy (amistre64):

http://mathforum.org/library/drmath/view/56863.html this has what i was trying to recall

OpenStudy (amistre64):

yes, but in a certain way

OpenStudy (amistre64):

fromthat site: Sometimes they may ask for two mean proportionals between a pair of numbers, and that is a lot tougher to work all the way through. Let's do an example to show what it means. Find two mean proportionals between 3 and 24. This means find two numbers, call them x and y, so that: 3 x y --- = --- = --- x y 24 If you work with this a little you can find that: 24x = y^2 and x^2 = 3y

OpenStudy (amistre64):

so, lets set that up :) because i apparently need the practice too 48 x y --- = --- = --- x y -750 is what we are tryng to determine

OpenStudy (anonymous):

48x-750=-36000 sqrt-36000=60i sqrt10

OpenStudy (amistre64):

i read there that it is setting up a common ratio, a geometric sequence

OpenStudy (amistre64):

lets try this way, since im more familiar with sequences than means an = a1 r^(n-1) a1 = 48 and a4=-750 lets find r a4 = 48 r^(3) -740/48 = r^3 cbrt(-740/48) = r now we can find any term in the sequence

OpenStudy (amistre64):

a2 = 48 cbrt(-740/48) = -119.46 a3 = 48 (cbrt(-740/48))^2 my calculator hates me at the moment lol

OpenStudy (anonymous):

So are those the answers? lol

OpenStudy (anonymous):

This is confusing me lol

OpenStudy (amistre64):

i think i had an error in it, but your original post looks good

OpenStudy (amistre64):

if each term is *2.5 from the next, itll work fine

OpenStudy (anonymous):

So, I was right?

OpenStudy (amistre64):

you was right

OpenStudy (anonymous):

Thank you for your help :)

OpenStudy (amistre64):

good luck